OpenStudy (anonymous):

Input 3 5 7 9 Output 11 17 23 29 Which function corresponds to the table of inputs and outputs? f(x) = 3x + 1 f(x) = 3x + 2 f(x) = x + 7 f(x) = 2x + 5

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

f(x) = 3x + 2 f(x) = 3(3) + 2 f(x) = 9 + 2 f(x) = 11 f(x) = 3x + 2 f(x) = 3(5) + 2 f(x) = 15 + 2 f(x) = 17 f(x) = 3x + 2 f(x) = 3(7) + 2 f(x) = 21 + 2 f(x) = 23 f(x) = 3x + 2 f(x) = 3(9) + 2 f(x) = 27 + 2 f(x) = 29